Please help me with this problem: A 5 mile taxi ride costs $5.25 and a 10 mi ride costs $8.75. Express the taxi cost as a function of the trip length. Then determine the cost of a 23 mi ride.
Answer by stanbon(57282)

Please help me with this problem:
A 5 mile taxi ride costs $5.25 and a 10 mi ride costs $8.75.
---
Two points: (5,5.25) (10,8.75)
------------------------------------
Express the taxi cost as a function of the trip length.
slope = (8.75-5.25)/(10-5) = 3.50/5 = 0.70
Intercept = ?
5.25 = 0.7*5 + b
b = 5.25-3.50
b = 1.75
----------------
C(x) = 0.7x+1.75
=======================
Then determine the cost of a 23 mi ride.
C(23) = 0.7*23+1.75
C(23) = $17.85
==================
Cheers,
Stan H.
==============